# Massively Parallel Minimum Spanning Tree in General Metric Spaces

### v1

@@ 3 Preliminaries
There are different regimes for the choice of $S$. We study the strictly sublinear regime (a.k.a. fully scalable algorithms). That is, given a graph of $n$ vertices, each machine has local space $n^{\delta}$, where $10<\delta<$ 1.
@@
    the last partition is groups all the vertices together, meaning $\mathcal{P}_{k}=\{V(G)\}$.
@@
Given a partition $\mathcal{P}$ and a set of edges $E$, we use $\mathcal{P}\oplus E$ to denote the partition obtained from joining the sets of $\mathcal{P}$ using the edges in $E$. That is, two vertices $s$ and $t$ are in the same set of $\mathcal{P}\oplus E$ if there exists a “path” of edges $(u_{0},v_{0}),(u_{1},v_{1}),\ldots,(u_{k},v_{k})$, where $s$ and $u_{0}$ are in the same of $\mathcal{P}$, $v_{i}$ and $u_{i+1}$ are in the same of $\mathcal{P}$ for all $i$, and $v_{k}$ and $t$ are in the same set of $\mathcal{P}$. Equivalently, one could define $\mathcal{P}\oplus E$ as the finest partition that is refined by both $\mathcal{P}$ and the set of connected components of $E$.
@@
Intuitively, starting with the empty graph, for each level $k$, the tree $T$ uses some edges $F_{\alpha^{k}}$ (i.e. $E_{\alpha^{k}}$) to join components of $\mathcal{P}_{\alpha^{k-1}}$, such that all the vertices within a set of $\mathcal{P}_{\alpha^{k}}$ are connected.
@@
[…]
Proposition 3.3 ([MillerPX13]).
@@ 4.1 The Algorithm
First, we give a brief description of Algorithm 1. Given a complete weighted graph $G$ such that the weights satisfy the triangle inequality, let $\alpha=\Theta\left(\frac{\log^{2}n}{\varepsilon}\right)$. To begin, the algorithm creates a hierarchy. The levels in the hierarchies are always indexed by $\{1,\alpha,\alpha^{2},\ldots,W\}$. Ideally, the algorithm would have used $\mathcal{C}=\{\mathcal{C}_{\alpha^{k}}\}_{k\geq 0}$ as its hierarchy, where $\mathcal{C}_{\alpha^{k}}$ is the set of connected components in $G_{\leq\alpha^{k}}$. As that is not possible under the 1vs2-Cycle conjecture, the algorithm uses an approximation of $\mathcal{C}$.
Initially, a hierarchy $\mathcal{P}^{\textnormal{\tiny MPX}}$ is obtained by intersecting MPX decompositions with different parameters. This hierarchy satisfies the same diameter and crossing probability guarantees as in Proposition 3.3. Another hierarchy $\mathcal{P}$ is defined by extending $\mathcal{P}^{\textnormal{\tiny MPX}}$. For each level $t=\alpha^{k}$, the partition $\mathcal{P}_{t}$ is built by joining the sets in $\mathcal{P}^{\textnormal{\tiny MPX}}_{t}$ using edges of weight at most $t$ that do not cross $\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha t}$. In short, $\mathcal{P}_{t}$ is the set of connected components in $G_{\leq t}[\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha t}]$. This hierarchy is a suitable approximation of $\mathcal{C}$, however, it cannot be computed efficiently either. Therefore, the algorithm uses a hierarchy $\hat{\mathcal{P}}$ as an approximation to $\mathcal{P}$.
@@
8       Run $\Theta\left(\log\frac{1}{\varepsilon}+\log\log n\right)$ rounds of the modified Borůvka algorithm (Algorithm 2), with initial components $\widehat{\mathcal{P}}_{t/\alpha}$, and edges $G_{\leq\alpha t}[\widehat{\mathcal{P}}_{t}]$, call the output edges $E^{1}_{t}$.
@@
Now we give a more detailed description of the algorithm. The parameter $\alpha$ is chosen such that such that $\alpha\geq\frac{L\log n}{\varepsilon}$ where $L=\log_{\alpha}n$ is the number of levels, e.g. take $\alpha=\Theta\left(\frac{\log^{2}n}{\varepsilon}\right)$. Creating the MPX decomposition $\{\mathcal{P}^{0}_{t}\}_{t=\alpha^{k}}$ can be done independently for each level in $O(1)$ rounds, where each set in $\mathcal{P}^{0}_{t}$ has diameter $t$ with high probability, and an edge of weight $w$ crosses $\mathcal{P}^{0}_{t}$ with probability $O\left(\frac{w\log n}{t}\right)$ (Proposition 3.3). Note that the sequence MPX decompositions $\mathcal{P}^{0}=\{\mathcal{P}^{0}_{\alpha^{k}}\}_{k\geq 0}$ is not a hierarchy, which is necessary for our application.
The MPX hierarchy $\mathcal{P}^{\textnormal{\tiny MPX}}=\{\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha^{k}}\}_{k\geq 0}$ is built by intersecting the decompositions. Formally, we let two vertices be in the same set of $\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha^{k}}$, if and only if they are in the same set of $\mathcal{P}^{0}_{\alpha^{i}}$ for all $i\geq k$. Observe, that by intersecting the partitions of the next levels, we ensure that each partition is a refinement of the next. It can be seen that the diameter and the crossing probability guarantees still hold for $\mathcal{P}^{\textnormal{\tiny MPX}}$ (Claim 4.2).
@@
1 Input: A weighted graph $G$, an initial partition $\widehat{\mathcal{P}}_{t/\alpha}$, a set of edges $E$, and the number of rounds $r$.
@@
To finalize, the algorithm makes sure $\widehat{\mathcal{P}}_{t}=\widehat{\mathcal{P}}_{t/\alpha}\oplus E_{t}$ by arbitrarily connecting any remaining components within a set of $\widehat{\mathcal{P}}_{t}$. Formally, let $\mathfrak{C}$ be the set of components Algorithm 2 halts on. For each set of vertices $A\in\widehat{\mathcal{P}}_{t}$, let $\mathfrak{C}_{A}$ be the components of $\mathfrak{C}$ inside $A$ (note that $\mathfrak{C}$ is a refinement of $\widehat{\mathcal{P}}_{t}$). Take an arbitrary component $C^{*}\in\mathfrak{C}_{A}$, and for any other component $C\in\mathfrak{C}_{A}$, add one of the edges between $C^{*}$ and $C$ to $E^{2}_{t}$.
The remainder of this section is devoted to the proof of Theorem 1. First, we analyze the approximation ratio (Lemma 4.1). Then, we go over the runtime analysis and the implementation details (Lemma 4.8).
@@ 4.2 Approximation Ratio
Algorithm 1 computes a $(1+\varepsilon)$-approximation of the minimum spanning tree, in expectation.
@@
The sequence of partitions $\mathcal{P}^{\textnormal{\tiny MPX}}=\{\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha^{k}}\}_{k\geq 0}$ is a hierarchy, and the guarantees of Proposition 3.3 still hold. That is, each set in $\mathcal{P}^{\textnormal{\tiny MPX}}_{t}$ has diameter at most $t$ w.h.p. and any edge of weight $w$ crosses $\mathcal{P}^{\textnormal{\tiny MPX}}_{t}$ with probability at most $O\left(\frac{w\log n}{t}\right)$.
@@
The diameter of any set in $\mathcal{P}^{\textnormal{\tiny MPX}}_{t}$ is at most $t$ with high probability. Partition $\mathcal{P}^{\textnormal{\tiny MPX}}_{t}$ refines $\mathcal{P}^{(0)}_{t}$. Therefore, any set $A\in\mathcal{P}^{\textnormal{\tiny MPX}}_{t}$ is the subset of some set $B\in\mathcal{P}^{0}_{t}$. By Proposition 3.3, the diameter of $B$ is at most $t$ with high probability. Therefore, since $G$ is a complete graph that satisfies the triangle inequality, the diameter of $A\subseteq B$ is also at most $t$.
The crossing probability of an edge is at most doubled for $\mathcal{P}^{\textnormal{\tiny MPX}}_{t}$ due to taking intersections. Take an edge $(u,v)$. The probability that it crosses $\mathcal{P}^{0}_{\alpha^{i}}$ is at most $O\left(\frac{(\log n)w(u,v)}{\alpha^{i}}\right)$, by Proposition 3.3. Also, $(u,v)$ crosses $\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha^{k}}$ only if it crosses some $\mathcal{P}^{0}_{\alpha^{i}}$ for $i\geq k$. Therefore, by taking the union bound over $i$, the probability that $(u,v)$ crosses $\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha^{k}}$ is at most:
@@
Observe that $G_{\leq t}[\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha t}]$ is a subgraph of $G_{\leq\alpha t}[\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha^{2}t}]$. Therefore, $\mathcal{P}_{t}$ is a refinement of $\mathcal{P}_{\alpha t}$, and $\mathcal{P}$ is a hierarchy. Take a set $A\in\mathcal{P}_{t}$. By definition $A$ is a connected component of $G_{\leq t}[\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha t}]$. Because the edges of $G_{\leq t}[\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha t}]$ do not cross $\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha t}$, $A$ is fully contained in a set $B\in\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha t}.$ By Claim 4.2, $B$ has diameter at most $\alpha t$. Therefore $A$ also has diameter at most $\alpha t$. Finally, for an edge with weight $w\leq t$ to cross $\mathcal{P}_{t}$, it must also cross $\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha t}$, which happens with probability at most $O\left(\frac{w\log n}{\alpha t}\right)$ (Claim 4.2). ∎
@@
$\sum_{t}t\left(\lvert\mathcal{P}_{t/\alpha}\rvert-\lvert\mathcal{P}_{t}\rvert\right)\leq\sum_{t}t\left(\lvert\mathcal{P}_{t/\alpha}\rvert-\lvert\mathcal{C}_{t/\alpha}\rvert\right)+\sum_{t}t\left(\lvert\mathcal{C}_{t/\alpha}\rvert-\lvert\mathcal{C}_{t}\rvert\right).$ (2)
@@
The second sum on the right-hand side of (2) has already been bounded by $\alpha\cdot w(\textnormal{MST}(G))$ in (1). It remains to bound the first part. Let $\tilde{E}_{t}$ be the edges of MST with weight $w\in(t/\alpha,t]$, i.e. $\mathcal{C}_{t}=\mathcal{C}_{t/\alpha}\oplus\tilde{E}_{t}$. The difference in the number of components $\lvert\mathcal{P}_{t/\alpha}\rvert-\lvert\mathcal{C}_{t/\alpha}\rvert$ is equal to the number of edges of weight at most $t$ in the MST (i.e. the edges in $\bigcup_{t^{\prime}\leq t/\alpha}\tilde{E}_{t^{\prime}}$) that cross $\mathcal{P}_{t/\alpha}$. For an edge in $\tilde{E}_{t^{\prime}}$ the probability of crossing $\mathcal{P}_{t/\alpha}$ is at most $O\left(\frac{t^{\prime}\log n}{t}\right)$ (by Claim 4.3). Therefore, we have:
@@
$\operatorname{\textbf{E}}\left[\sum_{t}t\left(\lvert\mathcal{P}_{t/\alpha}\rvert-\lvert\mathcal{C}_{t/\alpha}\rvert\right)\right]$ $\leq\sum_{t^{\prime}}\lvert\tilde{E}_{t^{\prime}}\rvert\cdot O\left({t^{\prime}\log n}\right)\cdot L$
@@
$\leq O\left({L\log n}\right)\cdot\sum_{t^{\prime}}t^{\prime}(\lvert\mathcal{C}_{t^{\prime}/\alpha}\rvert-\lvert\mathcal{C}_{t^{\prime}}\rvert)$
@@
At level $t$, we “lift” all the edges of $\tilde{F}_{t}$ that have not been captured by $\mathcal{P}_{t}$ to the minimum weight of the next level, meaning we let $\tilde{w}(e)=t$ for such an edge $e$. These edges remain unaccounted for. The rest of the edges in $\tilde{F}_{t}$ will be accounted for by the time we are through with this level. We go over the unaccounted edges of $\tilde{F}_{t}$ that have been captured by $\mathcal{P}_{t}$, and add an edge to $T$ if doing so does not create a cycle. In this case, we charge $\tilde{w}(e)$ to $e$ itself. Note that $\tilde{w}(e)\geq w(e)$ and $e$ is accounted for. After going through these edges, if a set in $\mathcal{P}_{t}$ is not fully connected by $T$, we make it connected by adding edges of weight at most $t$, call these extra edges. This can be done since $\mathcal{P}_{t}$ is connected in $G_{\leq t}$ (by Claim 4.3). We let $E_{t}$ be the set of added edges for this level.
@@
To conclude the proof, note that any edge $e\in\tilde{T}$ is at some point captured and hence accounted for, because the last level of the hierarchy groups all the vertices together. Also, when a vertex $e\in\tilde{T}$ charges $\tilde{w}(e)$ to an edge $e^{\prime}\in T$, it holds that $\tilde{w}(e)\geq w(e^{\prime})$. Therefore, it holds that $w(T)\leq\tilde{w}(\tilde{T})$. It remains to analyze $\operatorname{\textbf{E}}[\tilde{w}(\tilde{T})]$. For an edge $e\in\tilde{T}$ to be lifted at level $t$, it must have $w(e)\leq t$ and it must cross $\mathcal{P}_{t}$ which happens with probability $O\left(\frac{w(e)\log n}{\alpha t}\right)$. In case it is lifted, $\tilde{w}(e)$ is increased by at most $t-t/\alpha\leq t$. Therefore, it holds that
@@
$\operatorname{\textbf{E}}[\tilde{w}(\tilde{T})]\leq(1+\varepsilon)w(\tilde{T}).$
Combining this with $w(T)\leq\tilde{w}(\tilde{T})$ gives the result:
@@
Let the hierarchy $\hat{\mathcal{P}}$ be obtained as in Algorithm 1. Then, it holds that
@@
$\sum_{t}t\left(\lvert\widehat{\mathcal{P}}_{t/\alpha}\rvert-\lvert\widehat{\mathcal{P}}_{t}\rvert\right)\leq\sum_{t}t\left(\lvert\mathcal{P}_{t}\rvert-\lvert\widehat{\mathcal{P}}_{t}\rvert\right)+\sum_{t}t\left(\lvert\mathcal{P}_{t/\alpha}\rvert-\lvert\mathcal{P}_{t}\rvert\right),$ (3)
where the inequality holds because the terms involving $\mathcal{P}_{t}$ cancel out, and $\lvert\widehat{\mathcal{P}}_{t/\alpha}\rvert\leq\lvert\mathcal{P}_{t/\alpha}\rvert$. The second sum has already been bounded in Claim 4.4. It remains to analyze the first.
@@
[…]
@@
[…]
@@
where the second inequality was proven in Claim 4.4. Choosing a large enough constant for $r=\Theta\left(\log\frac{1}{\varepsilon}+\log\log n\right)$, and plugging this back into (3), gives the claim. ∎
@@
Let $\hat{\mathcal{P}}$ be obtained as in Algorithm 1. Then it holds that:
@@
To analyze the weight of $T$, take the two kinds of edges we have added to it. The edges from $\tilde{T}$ that have total weight at most $w(\tilde{T})$, and the arbitrary edges used to connect the remaining components which have total weight at most $\sum_{t}\alpha t(\lvert\mathcal{P}_{t}\rvert-\lvert\widehat{\mathcal{P}}_{t}\rvert)$. Therefore, we have
@@
$\leq\operatorname{\textbf{E}}[w(\tilde{T})]+\operatorname{\textbf{E}}\left[\sum_{t}\alpha t(\lvert\mathcal{P}_{t}\rvert-\lvert\widehat{\mathcal{P}}_{t}\rvert)\right]$
@@
For a high-level proof of why $T=\cup_{k\geq 0}E_{\alpha_{k}}$ approximates the MST w.r.t. $(G_{\leq\alpha t}\mid\widehat{\mathcal{P}}_{t}$), observe that if the modified Borůvka’s algorithm were executed till completion instead of stopping after $r=\Theta\left(\log\frac{1}{\varepsilon}+\log\log n\right)$ rounds, then the exact MST w.r.t. $(G_{\leq\alpha t}\mid\widehat{\mathcal{P}}_{t}$) would have been recovered. However, running the subroutine for $r$ rounds guarantees that at most an $\textnormal{poly}\left(\frac{\varepsilon}{\log n}\right)$-fraction of the edges of the exact MST are “missed” because of the early termination. For the arbitrary edges added in their place, we incur a multiplicative error of $\textnormal{poly}(\alpha)$ which is dominated by $2^{\Omega(r)}\varepsilon$. Therefore, overall we get a $(1+\varepsilon)$-approximation of the MST. Below is the formal proof.
Proof of Lemma 4.1.

### v2

@@ 3 Preliminaries
There are different regimes for the choice of $S$. We study the strictly sublinear regime (a.k.a. fully scalable algorithms). That is, given a graph of $n$ vertices, each machine has local space $n^{\delta}$, where $0<\delta<1$.
@@
    the last partition groups all the vertices together, meaning $\mathcal{P}_{k}=\{V(G)\}$.
@@
Given a partition $\mathcal{P}$ and a set of edges $E$, we use $\mathcal{P}\oplus E$ to denote the partition obtained from joining the sets of $\mathcal{P}$ using the edges in $E$. That is, two vertices $s$ and $t$ are in the same set of $\mathcal{P}\oplus E$ if there exists a “path” of edges $(u_{0},v_{0}),(u_{1},v_{1}),\ldots,(u_{k},v_{k})\in E$, where $s$ and $u_{0}$ are in the same set of $\mathcal{P}$, $v_{i}$ and $u_{i+1}$ are in the same set of $\mathcal{P}$ for all $i$, and $v_{k}$ and $t$ are in the same set of $\mathcal{P}$. Equivalently, one could define $\mathcal{P}\oplus E$ as the finest partition that is refined by both $\mathcal{P}$ and the set of connected components of $E$.
@@
Intuitively, starting with the empty graph, for each level $k$, the tree $T$ uses some edges of $F_{\alpha^{k}}$ (i.e. $E_{\alpha^{k}}$) to join components of $\mathcal{P}_{\alpha^{k-1}}$, such that all the vertices within a set of $\mathcal{P}_{\alpha^{k}}$ are connected.
@@
A key ingredient of our algorithm is a low-diameter decomposition of graphs using random shifts due to [MPX13]. Given any weighted graph $G$, this is obtained by drawing, for each vertex $u$, a delay $\delta_{u}$ from the exponential distribution with expectation $\Theta(\frac{t}{\log n})$, i.e. the distribution with density function $f(x)=c\frac{\log n}{t}\exp\left(-c\frac{\log n}{t}x\right)$. Then, the center for $c(u)$ for a vertex $u$ is defined as the vertex that minimizes the shifted distance. That is, $c(u)=\operatorname*{arg\,min}_{v}d(u,v)-\delta(v)$, where $d(u,v)$ is the distance between $u$ and $v$ (which is equal to $w(u,v)$ in case of metrics). This results in a decomposition, hereafter referred to as an MPX decomposition with diameter $t$, comprised of sets of vertices with the same center. In their paper, they show:
Proposition 3.3 ([MPX13]).
@@ 4.1 The Algorithm
First, we give a brief description of Algorithm˜1. Given a complete weighted graph $G$ such that the weights satisfy the triangle inequality, let $\alpha=\Theta\left(\frac{\log^{2}n}{\varepsilon}\right)$. To begin, the algorithm creates a hierarchy. The levels in the hierarchies are always indexed by $\{1,\alpha,\alpha^{2},\ldots,W\}$. Ideally, the algorithm would have used $\mathcal{C}=\{\mathcal{C}_{\alpha^{k}}\}_{k\geq 0}$ as its hierarchy, where $\mathcal{C}_{\alpha^{k}}$ is the set of connected components in $G_{\leq\alpha^{k}}$. As that is not possible under the 1vs2-Cycle conjecture, the algorithm uses an approximation of $\mathcal{C}$.
Initially, a hierarchy $\mathcal{P}^{\textnormal{\tiny MPX}}$ is obtained by intersecting MPX decompositions with different parameters. This hierarchy satisfies the same diameter and crossing probability guarantees as in Proposition˜3.3. Another hierarchy $\mathcal{P}$ is defined by extending $\mathcal{P}^{\textnormal{\tiny MPX}}$. For each level $t=\alpha^{k}$, the partition $\mathcal{P}_{t}$ is built by joining the sets in $\mathcal{P}^{\textnormal{\tiny MPX}}_{t}$ using edges of weight at most $t$ that do not cross $\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha t}$. In short, $\mathcal{P}_{t}$ is the set of connected components in $G_{\leq t}[\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha t}]$. This hierarchy is a suitable approximation of $\mathcal{C}$; however, it cannot be computed efficiently either. Therefore, the algorithm uses a hierarchy $\hat{\mathcal{P}}$ as an approximation to $\mathcal{P}$.
@@
8   Run $\Theta\left(\log\frac{1}{\varepsilon}+\log\log n\right)$ rounds of the modified Borůvka algorithm (Algorithm˜2), with initial components $\widehat{\mathcal{P}}_{t/\alpha}$, and edges $G_{\leq\alpha t}[\widehat{\mathcal{P}}_{t}]$, call the output edges $E^{1}_{t}$.
@@
Now we give a more detailed description of the algorithm. The parameter $\alpha$ is chosen such that such that $\alpha\geq\frac{L\log n}{\varepsilon}$ where $L=\log_{\alpha}n$ is the number of levels, e.g. take $\alpha=\Theta\left(\frac{\log^{2}n}{\varepsilon}\right)$. Creating the MPX decomposition $\{\mathcal{P}^{0}_{t}\}_{t=\alpha^{k}}$ can be done independently for each level in $O(1)$ rounds, where each set in $\mathcal{P}^{0}_{t}$ has diameter $t$ with high probability, and an edge of weight $w$ crosses $\mathcal{P}^{0}_{t}$ with probability $O\left(\frac{w\log n}{t}\right)$ (Proposition˜3.3). Note that the sequence MPX decompositions $\mathcal{P}^{0}=\{\mathcal{P}^{0}_{\alpha^{k}}\}_{k\geq 0}$ is not a hierarchy, which is necessary for our application.
The MPX hierarchy $\mathcal{P}^{\textnormal{\tiny MPX}}=\{\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha^{k}}\}_{k\geq 0}$ is built by intersecting the decompositions. Formally, we let two vertices be in the same set of $\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha^{k}}$, if and only if they are in the same set of $\mathcal{P}^{0}_{\alpha^{i}}$ for all $i\geq k$. Observe, that by intersecting the partitions of the next levels, we ensure that each partition is a refinement of the next. It can be seen that the diameter and the crossing probability guarantees still hold for $\mathcal{P}^{\textnormal{\tiny MPX}}$ (Claim˜4.2).
@@
1Input: A weighted graph $G$, an initial partition $\widehat{\mathcal{P}}_{t/\alpha}$, a set of edges $E$, and the number of rounds $r$.
@@
To finalize, the algorithm makes sure $\widehat{\mathcal{P}}_{t}=\widehat{\mathcal{P}}_{t/\alpha}\oplus E_{t}$ by arbitrarily connecting any remaining components within a set of $\widehat{\mathcal{P}}_{t}$. Formally, let $\mathfrak{C}$ be the set of components Algorithm˜2 halts on. For each set of vertices $A\in\widehat{\mathcal{P}}_{t}$, let $\mathfrak{C}_{A}$ be the components of $\mathfrak{C}$ inside $A$ (note that $\mathfrak{C}$ is a refinement of $\widehat{\mathcal{P}}_{t}$). Take an arbitrary component $C^{*}\in\mathfrak{C}_{A}$, and for any other component $C\in\mathfrak{C}_{A}$, add one of the edges between $C^{*}$ and $C$ to $E^{2}_{t}$.
The remainder of this section is devoted to the proof of Theorem˜1. First, we analyze the approximation ratio (Lemma˜4.1). Then, we go over the runtime analysis and the implementation details (Lemma˜4.8).
@@ 4.2 Approximation Ratio
Algorithm˜1 computes a $(1+\varepsilon)$-approximation of the minimum spanning tree, in expectation.
@@
The sequence of partitions $\mathcal{P}^{\textnormal{\tiny MPX}}=\{\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha^{k}}\}_{k\geq 0}$ is a hierarchy, and the guarantees of Proposition˜3.3 still hold. That is, each set in $\mathcal{P}^{\textnormal{\tiny MPX}}_{t}$ has diameter at most $t$ w.h.p. and any edge of weight $w$ crosses $\mathcal{P}^{\textnormal{\tiny MPX}}_{t}$ with probability at most $O\left(\frac{w\log n}{t}\right)$.
@@
The diameter of any set in $\mathcal{P}^{\textnormal{\tiny MPX}}_{t}$ is at most $t$ with high probability. Partition $\mathcal{P}^{\textnormal{\tiny MPX}}_{t}$ refines $\mathcal{P}^{(0)}_{t}$. Therefore, any set $A\in\mathcal{P}^{\textnormal{\tiny MPX}}_{t}$ is the subset of some set $B\in\mathcal{P}^{0}_{t}$. By Proposition˜3.3, the diameter of $B$ is at most $t$ with high probability. Therefore, since $G$ is a complete graph that satisfies the triangle inequality, the diameter of $A\subseteq B$ is also at most $t$.
The crossing probability of an edge is at most doubled for $\mathcal{P}^{\textnormal{\tiny MPX}}_{t}$ due to taking intersections. Take an edge $(u,v)$. The probability that it crosses $\mathcal{P}^{0}_{\alpha^{i}}$ is at most $O\left(\frac{(\log n)w(u,v)}{\alpha^{i}}\right)$, by Proposition˜3.3. Also, $(u,v)$ crosses $\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha^{k}}$ only if it crosses some $\mathcal{P}^{0}_{\alpha^{i}}$ for $i\geq k$. Therefore, by taking the union bound over $i$, the probability that $(u,v)$ crosses $\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha^{k}}$ is at most:
@@
Observe that $G_{\leq t}[\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha t}]$ is a subgraph of $G_{\leq\alpha t}[\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha^{2}t}]$. Therefore, $\mathcal{P}_{t}$ is a refinement of $\mathcal{P}_{\alpha t}$, and $\mathcal{P}$ is a hierarchy. Take a set $A\in\mathcal{P}_{t}$. By definition $A$ is a connected component of $G_{\leq t}[\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha t}]$. Because the edges of $G_{\leq t}[\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha t}]$ do not cross $\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha t}$, $A$ is fully contained in a set $B\in\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha t}.$ By Claim˜4.2, $B$ has diameter at most $\alpha t$. Therefore $A$ also has diameter at most $\alpha t$. Finally, for an edge with weight $w\leq t$ to cross $\mathcal{P}_{t}$, it must also cross $\mathcal{P}^{\textnormal{\tiny MPX}}_{\alpha t}$, which happens with probability at most $O\left(\frac{w\log n}{\alpha t}\right)$ (Claim˜4.2). ∎
@@
[…]
@@
The second sum on the right-hand side of (2) has already been bounded by $\alpha\cdot w(\textnormal{MST}(G))$ in (1). It remains to bound the first part. Let $\tilde{E}_{t}$ be the edges of MST with weight $w\in(t/\alpha,t]$, i.e. $\mathcal{C}_{t}=\mathcal{C}_{t/\alpha}\oplus\tilde{E}_{t}$. The difference in the number of components $\lvert\mathcal{P}_{t/\alpha}\rvert-\lvert\mathcal{C}_{t/\alpha}\rvert$ is equal to the number of edges of weight at most $t$ in the MST (i.e. the edges in $\bigcup_{t^{\prime}\leq t/\alpha}\tilde{E}_{t^{\prime}}$) that cross $\mathcal{P}_{t/\alpha}$. For an edge in $\tilde{E}_{t^{\prime}}$ the probability of crossing $\mathcal{P}_{t/\alpha}$ is at most $O\left(\frac{t^{\prime}\log n}{t}\right)$ (by Claim˜4.3). Therefore, we have:
@@
[…]
@@
[…]
@@
At level $t$, we “lift” all the edges of $\tilde{F}_{t}$ that have not been captured by $\mathcal{P}_{t}$ to the minimum weight of the next level, meaning we let $\tilde{w}(e)=t$ for such an edge $e$. These edges remain unaccounted for. The rest of the edges in $\tilde{F}_{t}$ will be accounted for by the time we are through with this level. We go over the unaccounted edges of $\tilde{F}_{t}$ that have been captured by $\mathcal{P}_{t}$, and add an edge to $T$ if doing so does not create a cycle. In this case, we charge $\tilde{w}(e)$ to $e$ itself. Note that $\tilde{w}(e)\geq w(e)$ and $e$ is accounted for. After going through these edges, if a set in $\mathcal{P}_{t}$ is not fully connected by $T$, we make it connected by adding edges of weight at most $t$, call these extra edges. This can be done since $\mathcal{P}_{t}$ is connected in $G_{\leq t}$ (by Claim˜4.3). We let $E_{t}$ be the set of added edges for this level.
@@
[…]
@@
[…]
@@
Let the hierarchy $\hat{\mathcal{P}}$ be obtained as in Algorithm˜1. Then, it holds that
@@
[…]
where the inequality holds because the terms involving $\mathcal{P}_{t}$ cancel out, and $\lvert\widehat{\mathcal{P}}_{t/\alpha}\rvert\leq\lvert\mathcal{P}_{t/\alpha}\rvert$. The second sum has already been bounded in Claim˜4.4. It remains to analyze the first.
@@
$\operatorname*{\textnormal{Pr}}\left(N^{(r)}_{A}\neq 1\right)=\operatorname*{\textnormal{Pr}}\left(N^{(r)}_{A}>1\right)\leq\operatorname{\textbf{E}}\left[N_{A}^{(r)}-1\right].$
@@
$\operatorname{\textbf{E}}[\lvert\mathcal{P}_{t}\rvert-\lvert\widehat{\mathcal{P}}_{t}\rvert$] $\leq\sum_{A\in\mathcal{P}_{t}}\operatorname*{\textnormal{Pr}}\left(N^{(r)}_{A}>1\right)$
@@
where the second inequality was proven in Claim˜4.4. Choosing a large enough constant for $r=\Theta\left(\log\frac{1}{\varepsilon}+\log\log n\right)$, and plugging this back into (3), gives the claim. ∎
@@
Let $\hat{\mathcal{P}}$ be obtained as in Algorithm˜1. Then it holds that:
@@
[…]
@@
[…]
@@
For a high-level proof of why $T=\cup_{k\geq 0}E_{\alpha_{k}}$ approximates the MST w.r.t. $(G_{\leq\alpha t}\mid\widehat{\mathcal{P}}_{t}$), observe that if the modified Borůvka’s algorithm were executed till completion instead of stopping after $r=\Theta\left(\log\frac{1}{\varepsilon}+\log\log n\right)$ rounds, then the exact MST w.r.t. $(G_{\leq\alpha t}\mid\widehat{\mathcal{P}}_{t}$) would have been recovered. However, running the subroutine for $r$ rounds guarantees that at most a $\textnormal{poly}\left(\frac{\varepsilon}{\log n}\right)$-fraction of the edges of the exact MST are “missed” because of the early termination. For the arbitrary edges added in their place, we incur a multiplicative error of $\textnormal{poly}(\alpha)$ which is dominated by $2^{\Omega(r)}\varepsilon$. Therefore, overall we get a $(1+\varepsilon)$-approximation of the MST. Below is the formal proof.
Proof of Lemma˜4.1.